# Adapt and Diffuse: Sample-adaptive Reconstruction via Latent Diffusion Models

### v2

@@ Abstract
Inverse problems arise in a multitude of applications, where the goal is to recover a clean signal from noisy and possibly (non)linear observations. The difficulty of a reconstruction problem depends on multiple factors, such as the structure of the ground truth signal, the severity of the degradation and the complex interactions between the above. This results in natural sample-by-sample variation in the difficulty of a reconstruction task, which is often overlooked by contemporary techniques. Our key observation is that most existing inverse problem solvers lack the ability to adapt their compute power to the difficulty of the reconstruction task, resulting in subpar performance and wasteful resource allocation. We propose a novel method that we call severity encoding, to estimate the degradation severity of noisy, degraded signals in the latent space of an autoencoder. We show that the estimated severity has strong correlation with the true corruption level and can give useful hints at the difficulty of reconstruction problems on a sample-by-sample basis. Furthermore, we propose a reconstruction method based on latent diffusion models that leverages the predicted degradation severities to fine-tune the reverse diffusion sampling trajectory and thus achieve sample-adaptive inference times. Our framework acts as a wrapper that can be combined with any latent diffusion-based baseline solver, imbuing it with sample-adaptivity and acceleration. We perform numerical experiments on both linear and nonlinear inverse problems and demonstrate that our technique greatly improves the performance of the baseline solver and achieves up to 10×10$\times$10 × acceleration in mean sampling speed.
@@ 1 Introduction
Inverse problems arise in a multitude of computer vision [26, 48], biomedical imaging [45] and scientific [14] applications, where the goal is to recover a clean signal from noisy and degraded observations. As information is fundamentally lost in the process, structural assumptions on the clean signal are needed to make recovery possible. Traditional compressed sensing [4] approaches utilize explicit regularizers that encourage sparsity of the reconstruction in transformation domains. More recently, data-driven deep learning methods have established new state-of-the-art in tackling most inverse problems (see an overview in [30]).
@@
Sample-adaptive methods that incorporate the difficulty of a reconstruction problem, or the severity of degradation, and allocate compute effort accordingly are thus highly desired. To the best of our knowledge, such methods have not been studied extensively in the literature. Unrolled networks [51, 46] have been proposed for reconstruction, that map the iterations of popular optimizers to learnable submodules, where deeper networks can be used to tackle more challenging reconstruction tasks. However, network size is determined in training time and therefore these methods are unable to adapt on a sample-by-sample basis. Deep Equilibrium Models have been proposed to solve inverse problems [13] by training networks of arbitrary depth through the construction of fixed-point iterations. These methods can adapt their compute in test time by scaling the number of iterations to convergence, however it is unclear how the optimal number of iterations correlates with degradation severity.
Diffusion models have established new state-of-the-art performance in synthesizing data of various modalities [11, 29, 31, 33, 36, 16, 35, 17, 25], inverse problem solving and image restoration [18, 37, 43, 8, 7, 5, 6, 23, 21, 22, 12]. Diffusion-based sampling techniques generate the missing information destroyed by the corruption step-by-step through a diffusion process that transforms pure noise into a target distribution. Recent work [7] has shown that the sampling trajectory can be significantly shortened by starting the reverse diffusion process from a good initial reconstruction, instead of pure noise. However, this approach treats the noise level of the starting manifold as a hyperparameter independent of degradation severity. Therefore, even though sampling is accelerated, the same number of function evaluations are required to reconstruct any sample.
More recently, latent domain diffusion, that is a diffusion process defined in the low-dimensional latent space of a pre-trained autoencoder, has demonstrated great success in image synthesis [33] and has been successfully applied to solving inverse problems [34, 39, 9] and in high-resolution image restoration [28]. Latent diffusion has the clear benefit of improved efficiency due to the reduced dimensionality of the problem leading to faster sampling. In addition to this, the latent space consists of compressed representations of relevant information in data and thus provides a natural space to quantify the loss of information due to image corruptions, which strongly correlates with the difficulty of the reconstruction task.
@@
Figure 1: Overview of our method: we estimate the degradation severity of corrupted images in the latent space of an autoencoder. We leverage the severity predictions to find the optimal start time in a latent reverse diffusion process on a sample-by-sample basis. As a result, inference cost is automatically scaled by the difficulty of the reconstruction task at test time.
@@ 2 Background
Diffusion models – Diffusion in the context of generative modeling consists of transforming a clean data distribution $\bm{x}_{0}\sim q_{0}(\bm{x})$ through a forward noising process, defined over $0\leq t\leq T,~{}t\in\mathbb{R}$, into some tractable distribution $q_{T}$. Typically, $q_{t}$ is chosen such that $\bm{x}_{t}$ is obtained from $\bm{x}_{0}$ via adding i.i.d. Gaussian noise, that is $q_{t}(\bm{x}_{t}|\bm{x}_{0})\sim\mathcal{N}(\bm{x}_{0},\sigma_{t}^{2}\mathbf{I})$, where $\sigma_{t}^{2}$ is from a known variance schedule. Diffusion models (DMs) [38, 15, 41, 42] learn to reverse the forward corruption process in order to generate data samples starting from a simple Gaussian distribution. The forward process can be described as an Itô stochastic differential equation (SDE) [44] $\text{d}\bm{x}=f(\bm{x},t)\text{d}t+g(t)\text{d}\bm{w},$ where $\bm{w}\in\mathbb{R}^{n}$ is the standard Wiener process. The forward SDE is reversible [1] and can be simulated given $\nabla_{\bm{x}}\log q_{t}(\bm{x})$, which is referred to as the score of the data distribution. The score is typically approximated by a neural network $s_{\bm{\theta}}(\bm{x}_{t},t)$ and learned from data. After discretizing the reverse SDE, a variety of discrete time sampling algorithms can be derived.
Denoising Diffusion Probabilistic Models (DDPMs) [38, 15] are obtained from the discretization of the variance preserving SDE with $f(\bm{x},t)=-\frac{1}{2}\beta_{t}\bm{x}$ and $g(t)=\sqrt{\beta_{t}}$, where $\beta_{t}$ is a pre-defined variance schedule that is a strictly increasing function of $t$. One can sample from the corresponding forward diffusion process at any time step $i$ as $\bm{x}_{i}=\sqrt{\bar{\alpha}_{i}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{i}}\bm{\varepsilon},$ with $\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{i},~{}\alpha_{i}=1-\beta_{i}$. By minimizing the denoising score-matching objective
@@
Latent Diffusion Models (LDMs) – LDMs [33] aim to mitigate the computational burden of traditional DMs by running diffusion in a low-dimensional latent space of an autoencoder. In particular, an encoder $\mathcal{E}_{0}$ is trained to extract a compressed representation $\bm{z}\in\mathbb{R}^{d},~{}d<<n$ of the input signal $\bm{x}$ in the form $\bm{z}=\mathcal{E}_{0}(\bm{x})$. To recover the clean signal from the latent representation $\bm{z}$, a decoder $\mathcal{D}_{0}$ is trained such that $\mathcal{D}_{0}(\mathcal{E}_{0}(\bm{x}))\approx\bm{x}$. A score model that progressively denoises $\bm{z}$ can be trained in the latent space of the autoencoder via the objective following the DDPM framework. The final generated image can be obtained by passing the denoised latent through $\mathcal{D}_{0}$.
@@
where the first term corresponds to an unconditional score model trained to predict noise on the signal without any information about the forward model $\mathcal{A}$. The score of the likelihood term however is challenging to estimate in general. Various approaches have emerged to incorporate the data acquisition model into a diffusion process, including projection-based approaches [43, 8, 7], restricting updates to stay on a given manifold [6], spectral approaches [21], or methods that tailor the diffusion process to the degradation [49, 12, 10]. [5] proposes Diffusion Posterior Sampling (DPS) that approximates the gradient of the likelihood as $\nabla_{\bm{x}_{i}}\log p(\bm{y}|\bm{x}_{i})\approx\nabla_{\bm{x}_{i}}\log p(\bm{y}|\hat{\bm{x}}_{0}(\bm{x}_{i}))$, where $\hat{\bm{x}}_{0}(\bm{x}_{i})$ is the posterior mean estimate obtained from the score model.
More recently, a flurry of activity has emerged to deploy LDMs for inverse problem solving. The most straightforward approach is to perform posterior sampling in latent space via the DPS approximation $\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i})))$ often referred to as Latent-DPS [39]. However, Latent-DPS achieves poor performance out-of-the-box due to the inaccuracy of the approximation in early stages of diffusion and due to additional complications introduced by the decoder [34, 39, 9]. To address the latter concern, [34] proposes two extensions of Latent-DPS: GML-DPS, that guides the diffusion towards fixed points of the encoder-decoder composition, and PSLD which adds a "gluing" term in order to avoid inconsistencies at mask boundaries. Authors in [39] take a different route and propose ReSample, a sampling technique that enforces data consistency directly on the posterior means by solving the optimization problem
@@
or its pixel domain equivalent, both referred to as hard data consistency. [9] leverages text-to-image LDMs and incorporates prompt-tuning into their algorithm while combining Latent-DPS and hard data consistency.
A key challenge of diffusion-based solvers is their heavy compute demand, as reconstructing a single sample requires typically $500-1000$ evaluations of a large score model. Come-Closer-Diffuse-Faster (CCDF) [7] shortens the sampling trajectory by leveraging a good initial posterior mean estimate $\hat{\bm{x}}_{0}$ from a reconstruction network. They initialize the reverse process by jumping to a fixed time step in the forward process via $\bm{x}_{k}=\sqrt{\bar{\alpha}_{k}}\hat{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{k}}\bm{\varepsilon},$ and only perform $k<<N$ reverse diffusion steps, where $k$ is a fixed hyperparameter.
@@ 3.1 Severity encoding
What is often referred to as the manifold hypothesis [2] states that natural images live in a lower dimensional manifold embedded in $n$-dimensional pixel-space. This in turn implies that the information contained in an image can be represented by a low-dimensional latent vector that encapsulates the relevant features of the image. Autoencoders [24, 32] learn a latent representation from data by first summarizing the input image into a compressed latent vector $\bm{z}=\mathcal{E}_{0}(\bm{x})$ through an encoder. Then, the original image can be recovered from the latent via the decoder $\hat{\bm{x}}=\mathcal{D}_{0}(\bm{z})$ such that $\bm{x}\approx\hat{\bm{x}}$. As the latent space of autoencoders contains only the relevant information of data, it is a more natural space to quantify the loss of information due to the degradation than the image domain.
In particular, assume that we have access to the latent representation of clean images $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x}_{0}),~{}\bm{z}_{0}\in\mathbb{R}^{d}$, for instance from a pre-trained autoencoder. We propose a severity encoder ^$\hat{\mathcal{E}}_{\theta}$ that achieves two objectives simultaneously: (1) it can predict the latent representation of a clean image, given a noisy and degraded observation and (2) it can quantify the error in its own latent estimation. We denote
@@
with $\hat{\bm{z}}\in\mathbb{R}^{d}$ the estimate of $\bm{z}_{0}$ and $\hat{\sigma}\in\mathbb{R}$ the estimated degradation severity to be specified shortly. We use the notation ^$\hat{\mathcal{E}}_{\bm{z}}(\bm{y})=\hat{\bm{z}}$ and ^$\hat{\mathcal{E}}_{\sigma}(\bm{y})=\hat{\sigma}$ for the two conceptual components of our model, however in practice a single architecture is used to represent ^$\hat{\mathcal{E}}_{\theta}$. The first objective can be interpreted as image reconstruction in the latent space of the autoencoder: for $\bm{y}=\mathcal{A}(\bm{x})+\bm{n}$ and $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x})$, we have ^$\hat{\mathcal{E}}_{\bm{z}}(\bm{y})=\hat{\bm{z}}\approx\bm{z}_{0}$. The second objective captures the intuition that recovering $\bm{z}_{0}$ from $\bm{y}$ exactly may not be possible, and the prediction error is proportional to the loss of information about $\bm{x}$ due to the corruption. Thus, even though the predicted latent $\hat{\bm{z}}$ might be away from the true $\bm{z}_{0}$, the encoder quantifies the uncertainty in its own prediction. More specifically, we make the assumption that the prediction error in latent space can be modeled as zero-mean i.i.d. Gaussian. That is, we assume that
@@
with $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x}_{0})$ for a fixed, pre-trained encoder $\mathcal{E}_{0}$ and $\bar{\sigma}^{2}(\bm{y},\bm{z}_{0})=\frac{1}{d-1}\sum_{i=1}^{d}(\bm{e}^{(i)}-\frac{1}{d}\sum_{j=1}^{d}\bm{e}^{(j)})^{2}$ is the sample variance of the prediction error estimating $\sigma_{*}^{2}$. Here $\lambda_{\sigma}>0$ is a hyperparameter that balances between reconstruction accuracy ($L_{lat.rec.}$) and error prediction performance ($L_{err.}$). We empirically observe that even small loss values of $L_{lat.rec.}$ (that is fairly good latent reconstruction) may correspond to visible reconstruction error in image domain as semantically less meaningful features in image domain are often not captured in the latent representation. Therefore, we utilize an extra loss term that imposes image domain consistency with the ground truth image in the form

### v3

@@ Abstract
Inverse problems arise in a multitude of applications, where the goal is to recover a clean signal from noisy and possibly (non)linear observations. The difficulty of a reconstruction problem depends on multiple factors, such as the ground truth signal structure, the severity of the degradation and the complex interactions between the above. This results in natural sample-by-sample variation in the difficulty of a reconstruction problem. Our key observation is that most existing inverse problem solvers lack the ability to adapt their compute power to the difficulty of the reconstruction task, resulting in subpar performance and wasteful resource allocation. We propose a novel method, severity encoding, to estimate the degradation severity of corrupted signals in the latent space of an autoencoder. We show that the estimated severity has strong correlation with the true corruption level and can provide useful hints on the difficulty of reconstruction problems on a sample-by-sample basis. Furthermore, we propose a reconstruction method based on latent diffusion models that leverages the predicted degradation severities to fine-tune the reverse diffusion sampling trajectory and thus achieve sample-adaptive inference times. Our framework, Flash-Diffusion, acts as a wrapper that can be combined with any latent diffusion-based baseline solver, imbuing it with sample-adaptivity and acceleration. We perform experiments on both linear and nonlinear inverse problems and demonstrate that our technique greatly improves the performance of the baseline solver and achieves up to 10×10$\times$10 × acceleration in mean sampling speed [^1] Code is available at https://github.com/z-fabian/flash-diffusion.
Published at the 41st International Conference on Machine Learning, Vienna, Austria, 2024
@@ 1 Introduction
Inverse problems arise in a multitude of computer vision (Ledig et al., 2017; Wang et al., 2018), biomedical imaging (Sriram et al., 2020) and scientific (Hand et al., 2018) applications, where the goal is to recover a clean signal from noisy and degraded observations. As information is fundamentally lost in the process, structural assumptions on the clean signal are needed to make recovery possible. Traditional compressed sensing (Candes et al., 2006) approaches utilize explicit regularizers that encourage sparsity of the reconstruction in transformation domains. More recently, data-driven deep learning methods have established new state-of-the-art in tackling most inverse problems (see an overview in (Ongie et al., 2020)).
@@
Figure 1: Overview of our method: we estimate the degradation severity of corrupted images in the latent space of an autoencoder (Severity Encoder). We leverage the severity predictions ($\hat{\sigma}$) to find the optimal start time in a latent reverse diffusion process on a sample-by-sample basis. As a result, inference cost is automatically scaled by the difficulty of the reconstruction task at test time.
Sample-adaptive methods that incorporate the difficulty of a reconstruction problem, or the severity of degradation, and allocate compute effort accordingly are thus highly desired. To the best of our knowledge, such methods have not been studied extensively in the literature. Unrolled networks (Zhang & Ghanem, 2018; Sun et al., 2016) have been proposed for reconstruction, that map the iterations of popular optimizers to learnable submodules, where deeper networks can be used to tackle more challenging reconstruction tasks. However, network size is determined in training time and therefore these methods are unable to adapt on a sample-by-sample basis. Deep Equilibrium Models have been proposed to solve inverse problems (Gilton et al., 2021) by training networks of arbitrary depth through the construction of fixed-point iterations. These methods can adapt their compute in test time by scaling the number of iterations to convergence, however it is unclear how the optimal number of iterations correlates with degradation severity.
Diffusion models have established new state-of-the-art performance in synthesizing data of various modalities (Dhariwal & Nichol, 2021; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022b; Ho et al., 2022a; Saharia et al., 2022a; Ho et al., 2022b; Kong et al., 2020), inverse problem solving and image restoration (Kadkhodaie & Simoncelli, 2021; Saharia et al., 2021; Song et al., 2021b; Chung & Ye, 2022; Chung et al., 2022c, a, b; Kawar et al., 2021, 2022a, 2022b; Fabian et al., 2023). Diffusion-based sampling techniques generate the missing information destroyed by the corruption step-by-step through a diffusion process that transforms pure noise into a target distribution. Recent work (Chung et al., 2022c) has shown that the sampling trajectory can be significantly shortened by starting the reverse diffusion process from a good initial reconstruction, instead of pure noise. However, this approach treats the noise level of the starting manifold as a hyperparameter independent of degradation severity. Therefore, even though sampling is accelerated, the same number of function evaluations are required to reconstruct any sample.
More recently, latent domain diffusion, that is a diffusion process defined in the low-dimensional latent space of a pre-trained autoencoder, has demonstrated great success in image synthesis (Rombach et al., 2022) and has been successfully applied to solving inverse problems (Rout et al., 2024; Song et al., 2023; Chung et al., 2023) and in high-resolution (Luo et al., 2023) and real-world (Jiang et al., 2023) image restoration. Latent diffusion has the clear benefit of improved efficiency due to the reduced dimensionality of the problem leading to faster sampling. In addition to this, the latent space consists of compressed representations of relevant information in data and thus provides a natural space to quantify the loss of information due to image corruptions, which strongly correlates with the difficulty of the reconstruction task.
@@ 2 Background
Diffusion models – Diffusion in the context of generative modeling consists of transforming a clean data distribution $\bm{x}_{0}\sim q_{0}(\bm{x})$ through a forward noising process, defined over $0\leq t\leq T,~{}t\in\mathbb{R}$, into some tractable distribution $q_{T}$. Typically, $q_{t}$ is chosen such that $\bm{x}_{t}$ is obtained from $\bm{x}_{0}$ via adding i.i.d. Gaussian noise, that is $q_{t}(\bm{x}_{t}|\bm{x}_{0})\sim\mathcal{N}(\bm{x}_{0},\sigma_{t}^{2}\mathbf{I})$, where $\sigma_{t}^{2}$ is from a known variance schedule. Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2020a, b) learn to reverse the forward corruption process in order to generate data samples starting from a simple Gaussian distribution. The forward process can be described as an Itô stochastic differential equation (SDE) (Song et al., 2020) $\text{d}\bm{x}=f(\bm{x},t)\text{d}t+g(t)\text{d}\bm{w},$ where $\bm{w}\in\mathbb{R}^{n}$ is the standard Wiener process. The forward SDE is reversible (Anderson, 1982) and can be simulated given $\nabla_{\bm{x}}\log q_{t}(\bm{x})$, which is referred to as the score of the data distribution. The score is typically approximated by a neural network $s_{\bm{\theta}}(\bm{x}_{t},t)$ and learned from data. After discretizing the reverse SDE, a variety of discrete time sampling algorithms can be derived.
Denoising Diffusion Probabilistic Models (DDPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) are obtained from the discretization of the variance preserving SDE with $f(\bm{x},t)=-\frac{1}{2}\beta_{t}\bm{x}$ and $g(t)=\sqrt{\beta_{t}}$, where $\beta_{t}$ is a pre-defined variance schedule that is a strictly increasing function of $t$. One can sample from the corresponding forward diffusion process at any time step $i$ as $\bm{x}_{i}=\sqrt{\bar{\alpha}_{i}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{i}}\bm{\varepsilon},$ with $\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{i},~{}\alpha_{i}=1-\beta_{i}$. By minimizing the denoising score-matching objective
@@
Latent Diffusion Models (LDMs) – LDMs (Rombach et al., 2022) aim to mitigate the computational burden of traditional DMs by running diffusion in a low-dimensional latent space of an autoencoder. In particular, an encoder $\mathcal{E}_{0}$ is trained to extract a compressed representation $\bm{z}\in\mathbb{R}^{d},~{}d<<n$ of the input signal $\bm{x}$ in the form $\bm{z}=\mathcal{E}_{0}(\bm{x})$. To recover the clean signal from the latent representation $\bm{z}$, a decoder $\mathcal{D}_{0}$ is trained such that $\mathcal{D}_{0}(\mathcal{E}_{0}(\bm{x}))\approx\bm{x}$. A score model that progressively denoises $\bm{z}$ can be trained in the latent space of the autoencoder via the objective following the DDPM framework. The final generated image can be obtained by passing the denoised latent through $\mathcal{D}_{0}$.
@@
where the first term corresponds to an unconditional score model trained to predict noise on the signal without any information about the forward model $\mathcal{A}$. The score of the likelihood term however is challenging to estimate in general. Various approaches have emerged to incorporate the data acquisition model into a diffusion process, including projection-based approaches (Song et al., 2021b; Chung & Ye, 2022; Chung et al., 2022c), restricting updates to stay on a given manifold (Chung et al., 2022b), spectral approaches (Kawar et al., 2022a), or methods that tailor the diffusion process to the degradation (Welker et al., 2022; Fabian et al., 2023; Delbracio & Milanfar, 2023). Chung et al. (2022a) proposes Diffusion Posterior Sampling (DPS) that approximates the gradient of the likelihood as $\nabla_{\bm{x}_{i}}\log p(\bm{y}|\bm{x}_{i})\approx\nabla_{\bm{x}_{i}}\log p(\bm{y}|\hat{\bm{x}}_{0}(\bm{x}_{i}))$, where $\hat{\bm{x}}_{0}(\bm{x}_{i})$ is the posterior mean estimate obtained from the score model.
More recently, a flurry of activity has emerged to deploy LDMs for inverse problem solving. The most straightforward approach is to perform posterior sampling in latent space via the DPS approximation
$\nabla_{\bm{z}_{i}}\log p(\bm{y}|\bm{z}_{i})\approx\nabla_{\bm{z}_{i}}\log p(\bm{y}|\mathcal{D}_{0}(\hat{\bm{z}}_{0}(\bm{z}_{i}))),$
which can be viewed as a "vanilla" extension of DPS to the latent domain (Rout et al., 2024) . We referred to this technique as Latent-DPS following Song et al. (2023). However, Latent-DPS achieves poor performance out-of-the-box due to the inaccuracy of the approximation in early stages of diffusion and due to additional complications introduced by the decoder (Rout et al., 2024; Song et al., 2023; Chung et al., 2023). To address the latter concern, Rout et al. (2024) proposes two extensions of Latent-DPS: GML-DPS, that guides the diffusion towards fixed points of the encoder-decoder composition, and PSLD which adds a "gluing" term in order to avoid inconsistencies at mask boundaries. Authors in Song et al. (2023) take a different route and propose ReSample, a sampling technique that enforces data consistency directly on the posterior means by solving the optimization problem
@@
or its pixel domain equivalent, both referred to as hard data consistency. Chung et al. (2023) leverages text-to-image LDMs and incorporates prompt-tuning into their algorithm while combining Latent-DPS and hard data consistency.
A key challenge of diffusion-based solvers is their heavy compute demand, as reconstructing a single sample requires typically $500-1000$ evaluations of a large score model. Come-Closer-Diffuse-Faster (CCDF) (Chung et al., 2022c) shortens the sampling trajectory by leveraging a good initial posterior mean estimate $\hat{\bm{x}}_{0}$ from a reconstruction network. They initialize the reverse process by jumping to a fixed time step in the forward process via $\bm{x}_{k}=\sqrt{\bar{\alpha}_{k}}\hat{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{k}}\bm{\varepsilon},$ and only perform $k<<N$ reverse diffusion steps, where $k$ is a fixed hyperparameter. More recently, Rout et al. (2023) proposes leveraging an efficient second-order approximation for estimating the posterior mean, notably accelerating sampling.
@@ 3.1 Severity encoding
What is often referred to as the manifold hypothesis (Bengio et al., 2013) states that natural images live in a lower dimensional manifold embedded in $n$-dimensional pixel-space. This in turn implies that the information contained in an image can be represented by a low-dimensional latent vector that encapsulates the relevant features of the image. Autoencoders (Kingma & Welling, 2013; Razavi et al., 2019) learn a latent representation from data by first summarizing the input image into a compressed latent vector $\bm{z}=\mathcal{E}_{0}(\bm{x})$ through an encoder. Then, the original image can be recovered from the latent via the decoder $\hat{\bm{x}}=\mathcal{D}_{0}(\bm{z})$ such that $\bm{x}\approx\hat{\bm{x}}$. As the latent space of autoencoders contains only the relevant information of data, it is a more natural space to quantify the loss of information due to the degradation than the image domain.
In particular, assume that we have access to the latent representation of clean images $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x}_{0}),~{}\bm{z}_{0}\in\mathbb{R}^{d}$, for instance from a pre-trained autoencoder. We propose a severity encoder ^$\hat{\mathcal{E}}_{\bm{\theta}}$ that achieves two objectives simultaneously: (1) it can predict the latent representation of a clean image, given a noisy and degraded observation and (2) it can quantify the error in its own latent estimation. We denote
@@
with $\hat{\bm{z}}\in\mathbb{R}^{d}$ the estimate of $\bm{z}_{0}$ and $\hat{\sigma}\in\mathbb{R}$ the estimated degradation severity to be specified shortly. We use the notation ^$\hat{\mathcal{E}}_{\bm{z}}(\bm{y})=\hat{\bm{z}}$ and ^$\hat{\mathcal{E}}_{\sigma}(\bm{y})=\hat{\sigma}$ for the two conceptual components of our model, however in practice a single architecture is used to represent ^$\hat{\mathcal{E}}_{\bm{\theta}}$. The first objective can be interpreted as image reconstruction in the latent space of the autoencoder: for $\bm{y}=\mathcal{A}(\bm{x})+\bm{n}$ and $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x})$, we have ^$\hat{\mathcal{E}}_{\bm{z}}(\bm{y})=\hat{\bm{z}}\approx\bm{z}_{0}$. The second objective captures the intuition that recovering $\bm{z}_{0}$ from $\bm{y}$ exactly may not be possible, and the prediction error is proportional to the loss of information about $\bm{x}$ due to the corruption. Thus, even though the predicted latent $\hat{\bm{z}}$ might be away from the true $\bm{z}_{0}$, the encoder quantifies the uncertainty in its own prediction. More specifically, we make the assumption that the prediction error in latent space can be modeled as zero-mean i.i.d. Gaussian. That is, we assume that
@@
with $\bm{z}_{0}=\mathcal{E}_{0}(\bm{x}_{0})$ for a fixed, pre-trained encoder $\mathcal{E}_{0}$ and
$\bar{\sigma}^{2}(\bm{y},\bm{z}_{0})=\frac{1}{d-1}\sum_{i=1}^{d}(\bm{e}^{(i)}-\frac{1}{d}\sum_{j=1}^{d}\bm{e}^{(j)})^{2}$
is the sample variance of the prediction error estimating $\sigma_{*}^{2}$. Here $\lambda_{\sigma}>0$ is a hyperparameter that balances between reconstruction accuracy ($L_{lat.rec.}$) and error prediction performance ($L_{err.}$). We empirically observe that even small loss values of $L_{lat.rec.}$ (that is fairly good latent reconstruction) may correspond to visible reconstruction error in image domain as semantically less meaningful features in image domain are often not captured in the latent representation. Furthermore, small errors in latent space can be severely amplified in image space after decoding, depending on the smoothness of the pretrained decoder. Therefore, we utilize an extra loss term that imposes image domain consistency with the ground truth image in the form